Circular No. 7507

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2000 T2 (KUSHIDA-MURAMATSU) S. Nakano, Sumoto, Japan, reports the recovery by T. Oribe of comet

P/1993 X1 (= 1993t = 1993 XIX) on CCD frames obtained with the 1.03m reflector at Saji Observatory on Oct. 3 and 4. The images were clearly cometary with coma diameter 10′′. The indicated correction to the prediction by B. G. Marsden on MPC 31663 was ΔT = −0.04 day, but neither Nakano nor Marsden was able to obtain a link to the 1993–1995 data without dramatically systematic residuals, particularly in declination. Oribe later found faint images of the comet on frames obtained on Sept. 26. C. E. Delahodde, European Southern Observatory, independently recovered the comet (as a pointlike object measured by A. Maury) with the Danish 1.54m reflector on Oct. 8. A further orbit computation by Marsden indicated that it was possible to link the 2000 data to the observations made after the 1994 conjunction. This computation revealed that single-night candidates for the comet found by C. W. Hergenrother, Lunar and Planetary Laboratory, with the 2.3-m Steward Observatory reflector at Kitt Peak in Sept. and Nov. 1999 indeed belong to the comet, which showed a possible 5′′ tail in p.a. 270^{o} on the first occasion. Orbital elements satisfactorily linking the 49 observations during 1994 Dec. 8–1995 June 23 and 1999 Sept. 13– 2000 Oct. 8 are given on MPEC 2000-T45, together with the 1999–2000 observations.

2000 UT  α_{2000}  δ_{2000}  m_{1}
Oct. 3.71875  2^{h}48^{m}34.^{s}54  +13^{o}08′08′′.7  20.2

SUPERNOVA 2000ds IN NGC 2768 T. Puckett and G. Dowdle, Mountain Town, GA, report the discovery

of an apparent supernova (mag 17.9) on an unfiltered CCD frame (limiting mag 19.9) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on Oct. 10.40 UT. SN 2000ds is located at α = 9^{h}11^{m}36^{s}.24, δ = +60^{o}01′42′′.2 (equinox 2000.0), which is 9′′.1 west and 32′′.1 south of the center of NGC 2768. The new object was also present on an unfiltered CCD frame taken on Oct. 11.38, and it appeared at mag 18.1 on a frame taken on Oct. 12.34. SN 2000ds does not appear on unfiltered images taken by Puckett during 2000 Jan. 6–May 28, and it does not appear on Palomar Sky Survey images taken on 1996 Mar. 17 (limiting mag ∼ 21.0) and on 1995 Feb. 5 (limiting mag ∼ 21.0).

2000 October 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green